# Online-BLS: An Accurate and Efficient Online Broad Learning System for Data Stream Classification

### v1

@@ 1 Introduction
Online Learning (OL) Hoi et al. (2021) provides a crucial solution for learning knowledge from various real-world data streams, such as weather Elwell and Polikar (2011), call record Swithinbank et al. (1977), and stock market trading data Hu et al. (2015). Typical OL algorithms involve learning model weights from one by one instance Gunasekara et al. (2023). The generic OL pipeline is as follows: 1) The initial model comes with zero or random weights because no samples are received. 2) When a new sample arrives, its model weights are rapidly updated to provide a more accurate prediction for subsequent arrivals Yuan et al. (2022). Thus, a brilliant OL model must be efficient and accurate. While appealing, existing state-of-the-art OL models are generally optimized based on gradient descent algorithms Wen et al. (2024); Su et al. (2024). To be specific, when a new sample arrives, all of them execute an error backpropagation and weight update process. Thus, none of them can achieve an optimal solution for model weights.
@@
Recently, Broad Learning System (BLS) has been proposed as an efficient batch or incremental machine learning model Chen and Liu (2018); Chen et al. (2019); Gong et al. (2022). First, BLS utilizes sparse coding to map raw data to a low-dimensional feature space to form feature nodes. The enhancement nodes are then obtained by random orthogonal transformation and non-linear mapping of the feature nodes. Finally, feature and enhancement nodes are cascaded, and then the relation weights between them and ground truth are estimated by the ridge regression algorithm. Thus, its closed-form solutions and incremental learning capabilities shed light on tackling the above OL issue.
@@
To address the suboptimal model weight issue, it is urgent to develop a novel OL framework based on BLS with closed-form solutions. Fortunately, various data incremental BLS algorithms Chen and Liu (2018); Fu et al. (2022); Du et al. (2024); Zhong et al. (2024) have been proposed to enhance accuracy and robustness. By setting the number of training samples for each increment to 1, most of them can inherently tackle online machine learning tasks. Thanks to the outstanding generalization of BLS, most of them can also rigorously guarantee the optimality for each online model update. Despite the straightforwardness, they still face challenges in online machine learning tasks. The common and most urgent issue is that all of them are based on the matrix inverse operation, still with high computational complexity and low numerical stability. It means that many errors are incurred when solving for online model weights. In other words, their accuracy and efficiency need to be improved.
@@
In this paper, we develop a native Online-BLS framework to address the suboptimality issue of OL model weights. To our knowledge, this paper clarifies the above issue for the first time and succeeds in solving it using our Online-BLS framework. Its technical innovations are summarized as follows. First, we deduce an Effective Weight Estimation Algorithm (EWEA) based on Cholesky factorization and forward-backward substitution, which significantly improves the accuracy of Online-BLS and offers excellent theoretical error bounds. Second, an Efficient Online Updating Strategy (EOUS) is designed based on the rank-one update algorithm of the Cholesky factor Gill et al. (1974); SEEGER (2007) to improve the efficiency of Online-BLS. The time complexity analysis shows a significant reduction in time overhead. Furthermore, to demonstrate the superiority and flexibility of our Online-BLS framework, we propose a straightforward extension to achieve non-stationary data stream classification. Finally, the superiority and efficiency of Online-BLS are empirically confirmed by a comprehensive comparison with baselines.
@@ 2.1 Online Learning
OL is a family of machine learning algorithms for learning models incrementally from sequential data Hoi et al. (2021). Existing OL algorithms fall into two main categories. The first category refers to traditional OL algorithms modified from traditional machine learning models. Hoeffding Tree (HT) has been proposed to construct a decision tree online from data streams Hulten et al. (2001). Subsequently, an Adaptive Random Forest (ARF) has been developed to handle concept drift scenarios Gomes et al. (2017). Specifically, once concept drift is detected, ARF resets its base tree to adapt to the new concept. Moreover, traditional online machine learning methods also include Passive Aggressive (PA) Crammer et al. (2006), Perceptron Rosenblatt (1958), Leveraging Bagging (LB) Bifet et al. (2010), etc. Despite simplicity, all of them fail to extract effective features from streaming data and thus degrade their accuracy.
Inspired by deep learning Larochelle et al. (2009); Zhang et al. (2021), the second one is online deep learning, which aims to learn a deep neural network online. To be specific, Online Gradient Descent (OGD) has been proposed to update its weights with gradient when a sample arrives Zinkevich (2003). Given that network depth is difficult to determine in advance, Sahoo et al. (2018) proposed the Hedge Backpropagation (HBP) to dynamically integrate multiple layers of neural networks, which is of great significance for subsequent research. Subsequently, Ashfahani and Pratama (2019) proposed Autonomous Deep Learning (ADL) that can construct its network structure from scratch by adjusting its depth and width without an initial one. To solve the online hyperparameter optimization challenge of neural networks, Continuously Adaptive Neural network for Data streams (CAND) Gunasekara et al. (2022) has been proposed, which chooses the best model from a candidate pool of neural networks trained with different hyperparameters combination. Adaptive Tree-like Neural Network (ATNN) has been proposed to solve the catastrophic forgetting problem by choosing suitable positions on the backbone to grow branches for the new concept Wen et al. (2024). To address concept drift and sub-network optimization conflict issues, EODL has been proposed, including depth adaption and parameter adaption strategies Su et al. (2024). Despite significant progress, none of them avoids gradient descent. They execute an error backpropagation and weight update when a sample arrives. However, these algorithms are far from enough to ensure the optimality of their model weights in the subsequent sample prediction. Hence, developing an OL framework based on closed-form solutions becomes crucial.
@@ 2.2 Incremental Broad Learning System
Due to its efficiency and versatility, BLS has been widely used for a variety of tasks Lei et al. (2024); Gong et al. (2024); Guo et al. (2024); Yun et al. (2024); Zhang et al. (2024). In particular, substantial works have been presented for incremental machine learning tasks. The first data incremental algorithm for BLS is deduced from Greville’s theory Greville (1960), which we refer to I-BLS in this paper Chen and Liu (2018). However, the performance of I-BLS dramatically degrades when learning a new class or task. To this end, TiBLS and BLS-CIL have been proposed based on residual learning and graph regularization, respectively Fu et al. (2022); Du et al. (2024). Furthermore, motivated by the fact that I-BLS still fails to improve accuracy when the size of the training data is not roughly equal for each addition, Zhong et al. have proposed a Robust Incremental BLS (RI-BLS)  Zhong et al. (2024). Among them, TiBLS stacks a new BLS module on the current model whenever a new task arrives. It views each new training sample as a new task in an online machine learning scenario. Therefore, a new BLS model needs to be learned. Inevitably, it imposes an expensive computational overhead and does not meet the requirements for efficient OL. In contrast, the remaining three algorithms only need to fine-tune their weights based on new arrivals, which is relatively efficient. The details of these incremental BLS algorithms are presented in Appendix A. Despite their feasibility, none of them can avoid matrix inversion operations. Thus, the defects of matrix inversion come along with them.
@@ 3.2 Online-BLS Framework
where ||$\cdot$|| ⋅ | denotes the $L_{2}$ norm. Since equation 4 is a convex optimization problem, we derive its gradient with respect to $\mathbf{W}$ and set it to $\mathbf{0}$. Then, we have
@@ 3.4 Efficient Online Update Strategy
To avoid expensive Cholesky decompositions for each online update, we discuss below how to efficiently derive $\mathbf{L}^{(k)}$ from $\mathbf{L}^{(k-1)}$ and $\mathbf{a}_{k}$. Inspired by the rank-one update strategy SEEGER (2007), we first cascade the $k$-th broad feature $\mathbf{a}_{k}$ and the previous Cholesky factor $\mathbf{L}^{(k-1)}$ as $[\mathbf{a}_{k},\mathbf{L}^{(k-1)}]$. Undergoing a series of orthogonal Givens rotations, we have
@@ 4.1 Error Bound
The detailed proof can be referred to Higham (2002). Lemma 1 states that its solution satisfies a slightly perturbed system. Also, each entry in the perturbation matrix $\mathbf{\Gamma}_{1}$ is much smaller than the corresponding element in $\mathbf{L}$.
@@
[…]
@@
The proof also refers to  Higham (2002).
@@
$|\mathbf{B}-\mathbf{L}\mathbf{L}^{\top}\mathbf{\Delta}\mathbf{\hat{W}}|$ $=|\mathbf{\Gamma}||\mathbf{\Delta}\mathbf{\hat{W}}|$
@@
$|\mathbf{B}-\mathbf{L}\mathbf{L}^{\top}\mathbf{\Delta}\mathbf{\hat{W}}|$ $=|\mathbf{\Gamma}||\mathbf{\Delta}\mathbf{\hat{W}}|$
@@
$|\mathbf{B}-$ $(\mathbf{A}^{\top}\mathbf{A}+\lambda\mathbf{I})\mathbf{\Delta}\mathbf{\hat{W}}|$
@@ 5.1 Experiments on Stationary Datasets
The comparison methods include I-BLS Chen and Liu (2018), BLS-CIL Du et al. (2024), and RI-BLS Zhong et al. (2024). To ensure fairness, we set the parameters common to all comparison methods and Online-BLS to be the same. Specifically, $n_{1}$, $n_{2}$, $n_{3}$, and $n_{4}$ were set to $10$, $10$, $1000$, and $1$, respectively. The regularization parameter $\lambda$ was set to $81e-$ 8 for I-BLS, RI-BLS and Online-BLS. For BLS-CIL, $\lambda_{1}$ and $\lambda_{2}$ were set to $1$ in the original paper, which is unstable in the online learning task. Therefore, we tuned both to 0.1 to obtain better performance. To draw credible conclusions, we repeated each experiment 10 times by changing the order of streaming data and the initialization parameters.
@@ 5.4 Experiments with Concept Drift Dataset
In this section, the effectiveness and flexibility of our Online-BLS are validated in the concept drift scenario. To learn non-stationary data streams, we adopt a forgetting factor to force our Online-BLS to assign larger weights to new arrivals. Specifically, we simply set it to $0.99$ for all datasets. We compare our method with two classes of baselines used for data stream classification with concept drift. The first category is online machine learning algorithms, including HT Hulten et al. (2001) and ARF Gomes et al. (2017). The second one is SOTA online deep learning models such as HBP Sahoo et al. (2018), ADL Ashfahani and Pratama (2019), CAND Gunasekara et al. (2022), ATNN Wen et al. (2024), and EODL Su et al. (2024). Beyond OCA, four metrics are used. They are Balanced Accuracy (BACC), Average Balanced Accuracy (AVRBACC), F1 score (F1), and Matthews Correlation Coefficient (MCC). It is worth noting that all the metrics used in this paper, except OCE, are such that larger values represent better model performance. The details of these metrics can also be found in Appendix C. The parameter settings of our method are consistent with those in section 5.1. Tables 4, 5, 6, and 7 demonstrate the final results on the Hyperplane, SEA, Electricity, and CoverType datasets, respectively.
@@
[…]
@@
[…]
@@
[…]
@@ A.1 BLS and I-BLS
BLS Chen and Liu [2018] is essentially equivalent to a ridge regression model, except that it utilizes stochastic transformations and nonlinear mapping to extract nonlinear representations. Assume that the first input sample is $\mathbf{x}_{1}\in\mathbb{R}^{d}$, where $d$ is the number of feature dimensions. The $i$-th feature node group $\mathbf{z}_{i}\in\mathbb{R}^{n_{1}}$ is defined as:
@@
$\mathbf{F}=\begin{cases}\mathbf{Q}^{+}&{\text{if}\ \mathbf{Q}\neq\mathbf{0}}\\ \big{(}[\mathbf{a}_{1},\mathbf{a}_{2},\cdots,\mathbf{a}_{k-1}]^{\top}\big{)}^{+}\mathbf{D}(1+\mathbf{D}^{\top}\mathbf{D})^{-1}&{\text{if}\ \mathbf{Q}=\mathbf{0}}.\\ \end{cases}$
@@ A.2 RI-BLS
To remedy the drawbacks of long training time and poor accuracy of I-BLS, RI-BLS Zhong et al. [2024] has been proposed, which devises an update strategy for the weight matrix by recursively computing two memory matrices. During the initialization training phase (i.e., RI-BLS receives the first training sample), the estimated weight is
@@ A.3 BLS-CIL
Compared to RI-BLS, BLS-CIL Du et al. [2024] has been proposed, which has an additional class-correlation loss function and was initially designed to address class-incremental tasks. By adding a single training sample at a time, BLS-CIL can be well adapted to online classification tasks. For the initialization training phase, the weight matrix $\mathbf{W}^{(1)}$ can be expressed as $\mathbf{W}^{(1)}=\big{(}\mathbf{K}^{(1)}\big{)}^{-1}\mathbf{a}_{1}\mathbf{y}_{1}^{\top}$, where $\mathbf{K}^{(1)}=\mathbf{I}+\mathbf{a}_{1}\mathbf{a}_{1}^{\top}$. After BLS-CIL receives a new sample $\mathbf{x}_{k}$, its weight matrix $\mathbf{W}^{(k)}$ can be formalized as
@@
where $\mathbf{K}^{(k)}=\mathbf{K}^{(k-1)}+\mathbf{H}$, $\mathbf{H}=\mathbf{a}_{k}\mathbf{a}_{k}^{\top}+\lambda(\mathbf{a}_{k-1}\mathbf{a}_{k-1}^{\top}-\mathbf{a}_{k}\mathbf{a}_{k-1}^{\top}+\mathbf{a}_{k}\mathbf{a}_{k}^{\top}-\mathbf{a}_{k-1}\mathbf{a}_{k}^{\top})$, and $\lambda=\begin{cases}-\lambda_{1}&{\text{if}\ \mathbf{y}_{k}\neq\mathbf{y}_{k-1}}\\ +\lambda_{2}&{\text{if}\ \mathbf{y}_{k}=\mathbf{y}_{k-1}},\\ \end{cases}$ where $\lambda_{1}$ and $\lambda_{2}$ are positive hyper-parameters which are responsible for intra-class compact term and inter-class sparse term losses, respectively.

### v2

@@ 1 Introduction
Online Learning (OL) hoi2021online provides a crucial solution for learning knowledge from various real-world data streams, such as weather 5975223, call record swithinbank1977drift, and stock market trading data hu2015itrade. Typical OL algorithms involve learning model weights from one by one instance gunasekara2023survey. The generic OL pipeline is as follows: 1) The initial model comes with zero or random weights because no samples are received. 2) When a new sample arrives, its model weights are rapidly updated to provide a more accurate prediction for subsequent arrivals yuan2022recent. Thus, a brilliant OL model must be efficient and accurate. While appealing, existing state-of-the-art OL models are generally optimized based on gradient descent algorithms wen2024adaptive; su2024elastic. To be specific, when a new sample arrives, all of them execute an error backpropagation and weight update process. Thus, none of them can achieve an optimal solution for model weights.
@@
Recently, Broad Learning System (BLS) has been proposed as an efficient batch or incremental machine learning model 7987745; 8457525; 9380770. First, BLS utilizes sparse coding to map raw data to a low-dimensional feature space to form feature nodes. The enhancement nodes are then obtained by random orthogonal transformation and non-linear mapping of the feature nodes. Finally, feature and enhancement nodes are cascaded, and then the relation weights between them and ground truth are estimated by the ridge regression algorithm. Thus, its closed-form solutions and incremental learning capabilities shed light on tackling the above OL issue.
@@
To address the suboptimal model weight issue, it is urgent to develop a novel OL framework based on BLS with closed-form solutions. Fortunately, various data incremental BLS algorithms 7987745; fu2022task; 10086560; 10533441 have been proposed to enhance accuracy and robustness. By setting the number of training samples for each increment to 1, most of them can inherently tackle online machine learning tasks. Thanks to the outstanding generalization of BLS, most of them can also rigorously guarantee the optimality for each online model update. Despite the straightforwardness, they still face challenges in online machine learning tasks. The common and most urgent issue is that all of them are based on the matrix inverse operation, still with high computational complexity and low numerical stability. It means that many errors are incurred when solving for online model weights. In other words, their accuracy and efficiency need to be improved.
@@
In this paper, we develop a native Online-BLS framework to address the suboptimality issue of OL model weights. To our knowledge, this paper clarifies the above issue for the first time and succeeds in solving it using our Online-BLS framework. Its technical innovations are summarized as follows. First, we deduce an Effective Weight Estimation Algorithm (EWEA) based on Cholesky factorization and forward-backward substitution, which significantly improves the accuracy of Online-BLS and offers excellent theoretical error bounds. Second, an Efficient Online Updating Strategy (EOUS) is designed based on the rank-one update algorithm of the Cholesky factor gill1974methods; seeger2007low to improve the efficiency of Online-BLS. The time complexity analysis shows a significant reduction in time overhead. Furthermore, to demonstrate the superiority and flexibility of our Online-BLS framework, we propose a straightforward extension to achieve non-stationary data stream classification. Finally, the superiority and efficiency of Online-BLS are empirically confirmed by a comprehensive comparison with baselines.
@@ 2.1 Online Learning
OL is a family of machine learning algorithms for learning models incrementally from sequential data hoi2021online. Existing OL algorithms fall into two main categories. The first category refers to traditional OL algorithms modified from traditional machine learning models. Hoeffding Tree (HT) has been proposed to construct a decision tree online from data streams 10.1145/502512.502529. Subsequently, an Adaptive Random Forest (ARF) has been developed to handle concept drift scenarios gomes2017adaptive. Specifically, once concept drift is detected, ARF resets its base tree to adapt to the new concept. Moreover, traditional online machine learning methods also include Passive Aggressive (PA) crammer2006online, Perceptron rosenblatt1958perceptron, Leveraging Bagging (LB) bifet2010leveraging, etc. Despite simplicity, all of them fail to extract effective features from streaming data and thus degrade their accuracy.
Inspired by deep learning larochelle2009exploring; 9360872, the second one is online deep learning, which aims to learn a deep neural network online. To be specific, Online Gradient Descent (OGD) has been proposed to update its weights with gradient when a sample arrives zinkevich2003online. Given that network depth is difficult to determine in advance, sahoo2018online proposed the Hedge Backpropagation (HBP) to dynamically integrate multiple layers of neural networks, which is of great significance for subsequent research. Subsequently, ashfahani2019autonomous proposed Autonomous Deep Learning (ADL) that can construct its network structure from scratch by adjusting its depth and width without an initial one. To solve the online hyperparameter optimization challenge of neural networks, Continuously Adaptive Neural network for Data streams (CAND) gunasekara2022online has been proposed, which chooses the best model from a candidate pool of neural networks trained with different hyperparameters combination. Adaptive Tree-like Neural Network (ATNN) has been proposed to solve the catastrophic forgetting problem by choosing suitable positions on the backbone to grow branches for the new concept wen2024adaptive. To address concept drift and sub-network optimization conflict issues, EODL has been proposed, including depth adaption and parameter adaption strategies su2024elastic. Despite significant progress, none of them avoids gradient descent. They execute an error backpropagation and weight update when a sample arrives. However, these algorithms are far from enough to ensure the optimality of their model weights in the subsequent sample prediction. Hence, developing an OL framework based on closed-form solutions becomes crucial.
@@ 2.2 Incremental Broad Learning System
Due to its efficiency and versatility, BLS has been widely used for a variety of tasks 10536023; 10506974; 10565294; 10609504; 10130786. In particular, substantial works have been presented for incremental machine learning tasks. The first data incremental algorithm for BLS is deduced from Greville’s theory greville1960some, which we refer to I-BLS in this paper 7987745. However, the performance of I-BLS dramatically degrades when learning a new class or task. To this end, TiBLS and BLS-CIL have been proposed based on residual learning and graph regularization, respectively fu2022task; 10086560. Furthermore, motivated by the fact that I-BLS still fails to improve accuracy when the size of the training data is not roughly equal for each addition, Zhong et al. have proposed a Robust Incremental BLS (RI-BLS)  10533441. Among them, TiBLS stacks a new BLS module on the current model whenever a new task arrives. It views each new training sample as a new task in an online machine learning scenario. Therefore, a new BLS model needs to be learned. Inevitably, it imposes an expensive computational overhead and does not meet the requirements for efficient OL. In contrast, the remaining three algorithms only need to fine-tune their weights based on new arrivals, which is relatively efficient. The details of these incremental BLS algorithms are presented in Appendix A. Despite their feasibility, none of them can avoid matrix inversion operations. Thus, the defects of matrix inversion come along with them.
@@ 3.2 Online-BLS Framework
where $|\cdot|$ denotes the $L_{2}$ norm. Since equation 4 is a convex optimization problem, we derive its gradient with respect to $\mathbf{W}$ and set it to $\mathbf{0}$. Then, we have
@@ 3.4 Efficient Online Update Strategy
To avoid expensive Cholesky decompositions for each online update, we discuss below how to efficiently derive $\mathbf{L}^{(k)}$ from $\mathbf{L}^{(k-1)}$ and $\mathbf{a}_{k}$. Inspired by the rank-one update strategy seeger2007low, we first cascade the $k$-th broad feature $\mathbf{a}_{k}$ and the previous Cholesky factor $\mathbf{L}^{(k-1)}$ as $\mathbf{a}_{k},\mathbf{L}^{(k-1)}]$. Undergoing a series of orthogonal Givens rotations, we have
@@ 4.1 Error Bound
The detailed proof can be referred to higham2002accuracy. Lemma 1 states that its solution satisfies a slightly perturbed system. Also, each entry in the perturbation matrix $\mathbf{\Gamma}_{1}$ is much smaller than the corresponding element in $\mathbf{L}$.
@@
$u=\frac{1}{2}\times(\text{gap between}\,1\,\text{and next largest floating point number}).$
@@
The proof also refers to  higham2002accuracy.
@@
[…]
@@
[…]
@@
[…]
@@ 5.1 Experiments on Stationary Datasets
The comparison methods include I-BLS 7987745, BLS-CIL 10086560, and RI-BLS 10533441. To ensure fairness, we set the parameters common to all comparison methods and Online-BLS to be the same. Specifically, $n_{1}$, $n_{2}$, $n_{3}$, and $n_{4}$ were set to $10$, $10$, $1000$, and $1$, respectively. The regularization parameter $\lambda$ was set to 1e81e-8 for I-BLS, RI-BLS and Online-BLS. For BLS-CIL, $\lambda_{1}$ and $\lambda_{2}$ were set to $1$ in the original paper, which is unstable in the online learning task. Therefore, we tuned both to 0.1 to obtain better performance. To draw credible conclusions, we repeated each experiment 10 times by changing the order of streaming data and the initialization parameters.
@@ 5.4 Experiments with Concept Drift Dataset
In this section, the effectiveness and flexibility of our Online-BLS are validated in the concept drift scenario. To learn non-stationary data streams, we adopt a forgetting factor to force our Online-BLS to assign larger weights to new arrivals. Specifically, we simply set it to $0.99$ for all datasets. We compare our method with two classes of baselines used for data stream classification with concept drift. The first category is online machine learning algorithms, including HT 10.1145/502512.502529 and ARF gomes2017adaptive. The second one is SOTA online deep learning models such as HBP sahoo2018online, ADL ashfahani2019autonomous, CAND gunasekara2022online, ATNN wen2024adaptive, and EODL su2024elastic. Beyond OCA, four metrics are used. They are Balanced Accuracy (BACC), Average Balanced Accuracy (AVRBACC), F1 score (F1), and Matthews Correlation Coefficient (MCC). It is worth noting that all the metrics used in this paper, except OCE, are such that larger values represent better model performance. The details of these metrics can also be found in Appendix C. The parameter settings of our method are consistent with those in section 5.1. Tables 4, 5, 6, and 7 demonstrate the final results on the Hyperplane, SEA, Electricity, and CoverType datasets, respectively.
@@
Table 4: Experimental results ($\%$) on Hyperplane dataset. Note: Bold indicates the best result, and underlining denotes the second-best result.
@@
Table 5: Experimental results ($\%$) on SEA dataset. Note: Bold indicates the best result, and underlining denotes the second-best result.
@@
Table 6: Experimental results ($\%$) on Electricity dataset. Note: Bold indicates the best result, and underlining denotes the second-best result.
@@ A.1 BLS and I-BLS
BLS 7987745 is essentially equivalent to a ridge regression model, except that it utilizes stochastic transformations and nonlinear mapping to extract nonlinear representations. Assume that the first input sample is $\mathbf{x}_{1}\in\mathbb{R}^{d}$, where $d$ is the number of feature dimensions. The $i$-th feature node group $\mathbf{z}_{i}\in\mathbb{R}^{n_{1}}$ is defined as:
@@
[…]
@@ A.2 RI-BLS
To remedy the drawbacks of long training time and poor accuracy of I-BLS, RI-BLS 10533441 has been proposed, which devises an update strategy for the weight matrix by recursively computing two memory matrices. During the initialization training phase (i.e., RI-BLS receives the first training sample), the estimated weight is
@@ A.3 BLS-CIL
Compared to RI-BLS, BLS-CIL 10086560 has been proposed, which has an additional class-correlation loss function and was initially designed to address class-incremental tasks. By adding a single training sample at a time, BLS-CIL can be well adapted to online classification tasks. For the initialization training phase, the weight matrix $\mathbf{W}^{(1)}$ can be expressed as $\mathbf{W}^{(1)}=\big(\mathbf{K}^{(1)}\big)^{-1}\mathbf{a}_{1}\mathbf{y}_{1}^{\top}$, where $\mathbf{K}^{(1)}=\mathbf{I}+\mathbf{a}_{1}\mathbf{a}_{1}^{\top}$. After BLS-CIL receives a new sample $\mathbf{x}_{k}$, its weight matrix $\mathbf{W}^{(k)}$ can be formalized as
@@
[…]